The length of a rectangle is 3ft less than twice the width and the area of the rectangle is 14ft^2. Find the dimensions of the rectangle.
Answer by jorel1380(3719)

Let w be the width of the rectangle, and 2w-3 be the length. Then:
w(2w-3)=14
2w²-3w-14=0
(2w-7)(w+2)=0
w=7/2 or -2
The width is 3-1/2 ft, the length is 4 ft.
☺☺☺☺